OpenStudy (anonymous):

if 36^12-m=6^2m, what is the value of m?

OpenStudy (owlcoffee):

Try rewriting 36^12 as a base of 6: \[36^{12}=6^{2m} \iff (6^2)^{12}=6^{2m}\]

OpenStudy (anonymous):

Thank you!!!

OpenStudy (owlcoffee):

Glad to help.

OpenStudy (anonymous):

So m would equal 6?

OpenStudy (owlcoffee):

That's incorrect :(

OpenStudy (owlcoffee):

Try making use of the property: \[(A^m)^n=A^{m.n}\]

OpenStudy (anonymous):

This is what I have so far: (6^2)^(12-m)=6^2m 6^24-2m=6^2m 24-2m=2m 24=4m m=6 What am I doing wrong?

OpenStudy (owlcoffee):

Ah, that's a fault from my part, I did not see the "-m" there, Sorry. You are correct. Well done!

OpenStudy (anonymous):

Awesome! Thank you so much!